Oxygen gas reacts with powdered aluminum according to the reaction: 4Al(s)+3O2(g)→2Al2O3(s) What volume of O2 gas (in L), measured at 775 Torr and 23 ∘C , completely reacts with 52.1 g of Al?
Answer #1

Step 1 : Calculate the number of moles of Aluminium

Number of moles of Aluminium = mass of Aluminium/ Atomic mass of Aluminium

= 52.1/27

= 1.9296 moles.

Step 2 : Calculate the number of moles of Oxygen required.

According to the reaction

4 moles of Aluminium react with 3 moles of Oxygen, then

1.9296 moles of Aluminium can react with how many moles of oxygen.

Number of moles of oxygen = 1.9296 X 3 /4

=1.4412 moles.

Step 3: Calculate the volume of oxygen according to Ideal gas equation.

PV = nRT ( when the value of pressure is given in torr, then the value of universal gas constant (R) is 62.39 L.mm/k.mol.)

Temperature should be taken in Kelvin, 23 + 273 = 296 k.

V = nRT/P

= 1.4472 X 62.4 X 296 /775

= 34.49 L.